The HIV disease can be produced by three viruses (A, B, C). For mice, the probability of developing the disease if infected with the virus A is 1/3, with virus B is 2/3 and virus C is 1/7.

a) if a researcher inoculates a mouse with one of the viruses at random, what is the probability that the mouse develops the disease?

b) If a mouse is inoculated and develops the HIV disease, what is the probability that it was inoculated with virus C?
Answer #1

a) P(the mouse develops the disease) = P(virus A and disease) + P(virus B and disease) + P(virus C and disease)

= (1/3 x 1/3) + (1/3 x 2/3) + (1/3 x 1/7)

= 0.38095

b) Bayes' Theorem: P(A | B) = P(A & B)/P(B)

P(virus C | the mouse develops the disease) = P(virus C and disease) / P(the mouse develops the disease)

= (1/3 x 1/7)/0.38095

= 0.125